Robert has some nickels and dimes worth $1.85. He has 11 more dimes than nickels. How many of each type does he have?
Answer by mananth(16946)

Robert has some nickels and dimes worth $1.85. He has 11 more dimes than nickels. How many of each type does he have?
let number of nickels be x
and dimes be y
y=x+11

5x+10y = 185 (cents)
substitute y
5x +10(x+11) =185
15x +110=185
15x = 75
x =5 nickels
plug x in y=x+11